at the end of Year 2 was $146,000 and its at the end of Year 1 was $155,000. The $1,402,000 in Year 2. Cost of goods sold amounted to $821,000 in Year 2. The companys operating cycle for Year 2 Is closest to: (Round your to 1 declmal place.) 103.2 days
Answer #1

Operating cycle = Inventory period+Account receivable period

Account receivable turnover = Sales/Average receivable

Average receivable = (146000+155000/2) = 150500

Account receivable turnover = 1402000/150500 = 9.32 times

Days sales in receivable = 365/9.32 = 39.16 days

Inventory turnover = Cost of goods sold/Average inventory

Average inventory = (149000+141000/2) = 145000

Inventory turnover = 821000/145000 = 5.66 times

Days sales in inventory = 365/5.66 = 64.49 days

Operating cycle period = 39.16+64.49 = 103.65 days

So answer is c) 103.2 days